Just today, Fawlty Foods, Inc.'s common stock paid a $1.40 annual dividend per share and had a closing price of $21. Assume that the market's required return, or capitaliza¬tion rate, for this investment is 12 percent and that dividends are expected to grow at a constant rate forever. 


a. Calculate the implied growth rate in dividends. 

b. What is the expected dividend

 yield7 

c. What is the expected capital gains yield7. 
Answer #1

a.         P0 = D1/(ke – g) = [D0(1 + g)]/(ke – g)

$21 = [$1.40(1 + g)]/(0.12 – g)

$21(0.12 – g) = $1.40(1 + g)

$2.52 – $21(g) = $1.40 + $1.40(g)

$1.12 = $22.40(g)

g = $1.12/$22.40 = 0.05 or 5 percent

 

b.         Expected dividend yield = D1/P0 = D0(1 + g)/P0

     = $1.40(1 + 0.05)/$21 = $1.47/$21 = 0.07

 

c.         Expected capital gains yield = g = 0.05.


answered by: Muhamad Burhan